Thursday, August 9, 2018, 10:00 am EDT - Compton

From Hall A Wiki
Jump to: navigation, search

Back to Main Compton Page >> Compton Meetings >> Previous Meeting

1. Dial Toll-Free Number: 1-888-240-2560 (U.S. & Canada)
2. International participants dial:
Toll Number: 1-408-740-7256
3. Enter meeting ID:140287970

Desktop sharing: 


Meeting room=F224-225 in Cebaf Center


  • FADC firmware (firmware only option) and Fiber amp approved for purchse


  • Need to check mode-matching

Photon Detector

  • DAQ work
  • PMT/linearity update

Electron detector

  • EEL test-stand update
  • Manitoba update


Attending: Amali, Alexandre, Jinlong, Cip, Dalton, Juan Carlos, Kent

Excused: Dave, Bob, Adam

Laser status:

  • need measurements for mode match
  • Dave has tried to lock the cavity, but hasn't gotten a big enough dip to get the cavity lock
  • He is pretty convinced that the parameter space (steering) has been spanned, so mode matching is next thing to diagnose
  • (satellite beam coming off the amplifier. Probably not a problem, but kinda weird.)
  • Amali and Cip available to help check mode matching
  • Plan to do mode match work with Dave early next week
  • Python script through EPICS to run RWP scan. Need to implement RWP mover readback into this.
    • Cip and Amali working on that now. Expect to finish today.


  • Juan Carlos coming next week. Little bit of work. Won't get a chance for more serious work until september.
  • Jinlong also not around until early september.
  • try out higher frequency
  • planning to bring LED test appartus
  • has not yet run pedestal noise tests
  • also thinking of a overnight run
    • EPICS getting stuck, think this is fixed
    • also pedestal false asymmetry?
  • (Hall probably closes for good about mid-sept)
  • Some channels we didn't check. (BPMs?) Need to verify the secondary signals.
  • Check the GUIs, logging scripts, etc.

Electron Detector:

  • some runs taken with e-det test stand
  • small peak seen, but not clear it is real
  • looking for larger source
  • No Manitoba update


  • Lead-tungstate geometry to GSO geometry... but not working. Cip can help today.
  • Juan-Carlos will send the GSO gdml geometry that he was using